Freight car loadings over an 18-week period at a busy port are as follows:

Week Number Week Number Week Number
1 270 7 360 13 390
2 290 8 370 14 400
3 310 9 380 15 440
4 290 10 360 16 460
5 310 11 370 17 480
6 320 12 380 18 500


a.
Determine a linear trend line for expected freight car loadings. (Round your intermediate calculations and final answers to 2 decimal places.)

=___+___ t

b. Use the above trend equation to predict expected loadings for Weeks 20 & 21. (Round your final answers to 2 decimal places.)

The forecasted demand for Week 20 is ___ and for Week 21 is   ____

c. The manager intends to install new equipment when the volume exceeds 850 loadings per week. Assuming the current trend continues, in which week (at the earliest) should the loading volume reach that level? (Use the rounded answers, as required, from any previous part of this problem. Do not round any other intermediate calculations. Round your final answer to 2 decimal places.)

It should reach 850 loadings in Week ____
Answer #1


PERIOD

DEMAND

X

Y

XY

X^2

1

270

1

270

270

1

2

290

2

290

580

4

3

310

3

310

930

9

4

290

4

290

1160

16

5

310

5

310

1550

25

6

320

6

320

1920

36

7

360

7

360

2520

49

8

370

8

370

2960

64

9

380

9

380

3420

81

10

360

10

360

3600

100

11

370

11

370

4070

121

12

380

12

380

4560

144

13

390

13

390

5070

169

14

400

14

400

5600

196

15

440

15

440

6600

225

16

460

16

460

7360

256

17

480

17

480

8160

289

18

500

18

500

9000

324

SIGMA

171

6680

69330

2109


INTERCEPT(B0) = (SIGMA(Y) * SIGMA(X^2) - SIGMA(X) * SIGMA(XY)) / (N * SIGMA(X^2) - SIGMA(X)^2)

INTERCEPT = (6680 * 2109) - (171 * 69330) / ((18 * 2109) - 171^2) = 256.01

SLOPE(B1) = ((N * SIGMA(XY)) - (SIGMA(X) * SIGMA(Y))) - (N * SIGMA(X^2) - SIGMA(X)^2)

SLOPE = ((18 * 69330) - (171 * 6680) / ((18 * 2109) - 171^2) = 12.12

EQUATION

FORECAST = INTERCEPT + SLOPE(X)

FORECAST= 256.01 + (12.12 * (X))

FORECAST

1

270

268.13

2

290

280.25

3

310

292.37

4

290

304.49

5

310

316.61

6

320

328.73

7

360

340.85

8

370

352.97

9

380

365.09

10

360

377.21

11

370

389.33

12

380

401.45

13

390

413.57

14

400

425.69

15

440

437.81

16

460

449.93

17

480

462.05

18

500

474.17

19

486.29

20

498.41

21

510.53


Using line equation Y = A + B(x)

We can calculate x = (Y - A) / B

= (850 - 256.01) / 12.12 = 49.00907591 or 49 weeks

**PLEASE LEAVE A THUMBS UP :)